if v=4+2f, solve 3v+2f=28 for f
Answer by schrammbledeggs(41)

Substitution is the key to solving this problem.
We're given V = 4 + 2F and are asked to solve 3V + 2F = 28
So what we'll do is plug the value of V into the equation.
3 (4 + 2F) + 2F = 28
Now we distribute the 3.
(3 x 4) + (3 X 2F) + 2F = 28
12 + 6F + 2F = 28
Now we combine like terms (the F's)
12 + 8F = 28
Subtract 12 from both sides
12 + 8F - 12 = 28 - 12
8F = 16
Divide both sides by 8
8F/8 = 16/8
F = 2
:)